OpenStudy (anonymous):

Find the slope of the line whose equation is 8y = 4 - 3x. as an answer i got -3/8?

OpenStudy (anonymous):

you got it! divide both sides of the equation by 8 to get y = 4/8 - 3/8*x or y = 1/2 - 3/8*x

OpenStudy (anonymous):

correct again as slope is the same as coefficient of the x term